Opened 17 months ago

Closed 7 days ago

#607 closed flight processing (fixed)

EUFAR15/28, flight day 271b/2015, Mont Blanc

Reported by: asm Owned by:
Priority: immediate Milestone:
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by dac)

Data location: ~/arsf_data/2015/flight_data/france/EUFAR15_28-2015_271b_Mont_Blanc

Data arrived from ARSF via SATA disk on 26/10/2015

Scientific objective: Hyperspectral airborne measurements over snow, glaciers and vegetation; intercomparison with ground-based
measurements and modelling.

Priority: Unknown

PI: Antoine Rabatel

EUFAR Project ID: ISOTHERM

Sensors

  • Fenix (requested)
  • Leica LIDAR (requested)
  • Leica FW LIDAR (not requested but flown anyway)
  • Owl (not requested but flown anyway)

Change History (33)

comment:1 Changed 17 months ago by asm

General processing
For general processing, will use Rinex data provided by Christian Vincent (already placed on basestation directory).

comment:2 Changed 15 months ago by dac

  • Description modified (diff)

From Gary 05/01/2016 "all LiDAR and RCD to be processed"

comment:3 Changed 15 months ago by dac

Fenix

Replaced wavelengths and ran through radiometric calibration using aplcal to get a better idea of data quality.

Last edited 15 months ago by dac (previous) (diff)

comment:4 Changed 13 months ago by asm

Navigation Process

Started.

comment:5 Changed 13 months ago by asm

Basestation verification:

Latitude 45 55 49.13361
Longitude 6 54 52.48238
Ell. Heightt 2133.858 metres
AntHgt 0.440m (LEIAS10, MeasDist 0.360m)

comment:6 Changed 13 months ago by asm

Navigation Process

Really good results, accuracy is one order of magnitude better than limits. Copying files across and navigation process will be finished.

comment:7 Changed 13 months ago by asm

RCD Processing

Started. Have already converted raw files to tif.

comment:8 Changed 13 months ago by asm

RCD Processing

There are two eventfiles and two PhotoId. The one that gives right values is 150928-053347.

comment:9 Changed 13 months ago by asm

RCD Processing

Flightline 009 only had one photo not totally covered by clouds. Everything else looks OK, no data quality remarks. Delivery and readme created. RCD is waiting delivery check.

comment:10 Changed 13 months ago by dac

RCD

Started delivery check.

comment:11 Changed 13 months ago by dac

RCD Delivery Check

Looks OK. However, images being removed due to cloud coverage isn't mentioned in readme. Once this has been changed ready to deliver.

comment:12 Changed 13 months ago by stgo

Lidar processing

Lines generally have the boresight pitch and roll. Only exception is 141626, will adjust manually for this line.

pitch -0.00226
roll -0.00101

comment:13 Changed 13 months ago by stgo

Lidar processing

141626 pitch and roll adjustments:

pitch -0.00081
roll -0.00246

processing full waveform data

comment:14 Changed 13 months ago by asm

RCD processing

Have added a note on readme about clouds and pictures not included.

comment:15 Changed 13 months ago by dac

RCD

Updated readme looks good. Ready to delivery.

comment:16 Changed 13 months ago by asm

Delivery

Camera delivered on 17/03/2016 via FTP (arsf20) and notification has been sent to PI.

comment:17 Changed 13 months ago by dac

Fenix Processing

Set SCT files generating over weekend ready for gej next week.

comment:18 Changed 13 months ago by stgo

Lidar processing

Created a delivery, ready for check.

comment:19 Changed 12 months ago by gej

Fenix Processing

Data is very cloudy, with not many usable areas
found SCT values of:

Flightline FENIX
1 1.00
2 1.00
3 1.00
4 1.00
5 0.97
6 1.00
7 1.00

Mapping files now

comment:20 Changed 12 months ago by gej

Fenix Processing

Delivery and Readme created

Ready for Delivery check

comment:21 Changed 12 months ago by lah

Lidar DC

  • Roll & pitch look fine
  • There's a bit of missed cloud in a couple of flight lines that you might want to reconsider. They could be patches of mountain, but the surrounding bits are classified as noise, so would make sense to get classify these as noise too.
  • Noisy should still be spelled without the e in the readme
  • Corrected sol file name
  • demcompare masked mean = 4.44
  • FW files look ok in waveviewer

Everything ok apart from recheck that classification and the readme typo

comment:22 Changed 12 months ago by dac

LiDAR

(taking over from stgo)

  • Corrected typo in reame
  • Redid classification - copying to FW and redoing ASCII

comment:23 Changed 12 months ago by dac

LiDAR

Copied classification to FW files, recreated ASCII and recalculated point cloud statistics for readme.

Marking as ready to deliver.

comment:24 Changed 12 months ago by dac

  • Description modified (diff)

Fenix

Checking over notes from Gary after checking screenshots and level1b data PI requested Fenix data to be processed. This was reflected in quote agreed by Gary 26/02/2016

comment:25 Changed 12 months ago by asm

Fenix Delivery Check

Started.

comment:26 Changed 12 months ago by asm

Fenix Delivery Check

-Readme and logsheet says that there are only 6 flightlines but 7 flightlines have been included on this delivery. It is the first flight-line the one not included on logsheet we concluded it should be included on the delivery as it has usable data and a great overlapping region. Will need to update the readme to include that flightline and the correct table should look:

Logsheet flightline name Fenix file identifier
not on logsheet f271b011b
013 f271b021b
... ...

-Data quality remarks only says: There is a lot of cloud cover across all flightlines. Would be better to specify that caused the fenix saturation.

comment:27 Changed 12 months ago by dac

Fenix

Updated names in readme and made note "The first file was not on the logsheet (likely a test line) but has been included as it overlaps the other lines for the flight and may contain useful information."

Added "Due to this cloud cover there are a large number of bands with overflows" - Same as 271a.

Ready to go (once mapped files have been zipped) if Aser is happy comments from DC have been addressed.

comment:28 Changed 12 months ago by asm

Fenix Delivery Check

-Also one flightline 016 was aborted but could be delivered as it could have usable data.

-Apl commands tested and working fine.

-Py6S wrapper was failing to find igm files. Tried every single flightline separately and the spectra looks fine (could not find vegetation pixel for flightline 3 and 7)

-All others checks have been done and everything looks OK.

Will zip files files and will ready to go.

comment:29 Changed 12 months ago by asm

Delivery

Fenix and LiDAR delivered on 29/03/2016 via FTP (arsf20) and notification has been sent to PI. This finalises this project.

comment:30 Changed 12 months ago by dac

Archiving

Started uploading processed data to NEODC

comment:31 Changed 12 months ago by dac

Archiving

Processed data uploaded to NEODC

comment:33 Changed 7 days ago by asm

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available from NEODC. Closing ticket.

Note: See TracTickets for help on using tickets.